OpenStudy (anonymous):

express as a single fraction the exact value of sin 75

OpenStudy (anonymous):

your job is to recognize \(75=45+30\) and then use the addition angle formula do you know it?

OpenStudy (anonymous):

\[\sin(\alpha +\beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\] replace \(\alpha\) by 45 and \(\beta\) by 30 and evalutate